# Global solutions to cross-diffusion systems with independent advections in one dimension

### 20 Mar 2026 v1

@@ 3.1. A priori estimates
First, we obtain a priori estimates for the solutions of (1.3). While they are quite standard comparing to the rest of the paper, they seem to be quite new in the fast-diffusion regime (particularly, the $L^{\infty}$ estimates). Moreover, one has to be careful as we work in the full range of exponents $\alpha\in(0,\infty)$ so for instance, when $\alpha\in(0,\frac{1}{3}]$, it is well-known that the energy $\int_{\mathbb{R}}s_{\varepsilon}^{\alpha}\mathop{}\!\mathrm{d}x$ is no longer integrable as well as the second moment $\int_{\mathbb{R}}s_{\varepsilon}\,|x|^{2}\mathop{}\!\mathrm{d}x$ which can be deduced by looking at the Barenblatt profile, see e.g. [14, eq. (2.4)].
@@
Let $u_{\varepsilon}$, $v_{\varepsilon}$ be the solution of (1.3). Then, for all $T>0$, the following sequences are bounded
@@
    $\{u_{\varepsilon}|x|^{2}\}$, $\{v_{\varepsilon}|x|^{2}\}$ (if $\alpha>\frac{1}{3}$) and $\{u_{\varepsilon}|x|\}$, $\{v_{\varepsilon}|x|\}$ (if $\alpha\in(0,\frac{1}{3}]$) in $L^{\infty}(0,T;L^{1}(\mathbb{R}))$,
@@
    $\{u_{\varepsilon}\log u_{\varepsilon}\,|x|\}$, $\{v_{\varepsilon}\log v_{\varepsilon}|x|\}$ (for $\alpha>\frac{1}{3}$) and $\{u_{\varepsilon}\log u_{\varepsilon}\,|x|^{\frac{1}{2}}\}$, $\{v_{\varepsilon}\log v_{\varepsilon}|x|^{\frac{1}{2}}\}$ in $L^{\infty}(0,T;L^{1}(\mathbb{R}))$ ,
@@
    $\{\partial_{t}u_{\varepsilon}\}$, $\{\partial_{t}v_{\varepsilon}\}$ is bounded in $L^{2}(0,T;H^{-1}(\mathbb{R}))$,
@@
We will need the following computation borrowed from [23, Prop. 2.1].
@@
$\int_{0}^{T}\int_{\mathbb{R}}s_{\varepsilon}^{2-\alpha}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}t\leq\|s^{0}\|_{L^{1}_{x}}\,\int_{0}^{T}\|s_{\varepsilon}^{1-\alpha}\|_{L^{\infty}(\mathbb{R})}\mathop{}\!\mathrm{d}t\leq\int_{0}^{T}\int_{\mathbb{R}}|\partial_{x}s_{\varepsilon}^{1-\alpha}|\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}t\leq$
@@
We will also need the following lemma proved in [12, Lemma D.2].
@@
We divide the proof into few steps. As a preliminary observation, we note that by the conservation of mass $M:=\int_{\mathbb{R}}s_{\varepsilon}(t,x)\mathop{}\!\mathrm{d}x=\int_{\mathbb{R}}s^{0}(x)\mathop{}\!\mathrm{d}x$, $\int_{\mathbb{R}}u_{\varepsilon}(t,x)\mathop{}\!\mathrm{d}x=\int_{\mathbb{R}}u^{0}(x)\mathop{}\!\mathrm{d}x$, $\int_{\mathbb{R}}v_{\varepsilon}(t,x)\mathop{}\!\mathrm{d}x=\int_{\mathbb{R}}v^{0}(x)\mathop{}\!\mathrm{d}x$. In particular, Assumption 1.1 implies that all the sequences $\{\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\}$, $\{\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]\}$, $\{\partial^{2}_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\}$, $\{\partial^{2}_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]\}$ are uniformly bounded in the space $L^{\infty}((0,T)\times\mathbb{R})$ and we denote the bound by $\|\partial_{x}\mathcal{V}^{1}\|_{L^{\infty}}$, $\|\partial_{x}\mathcal{V}^{2}\|_{L^{\infty}}$, $\|\partial^{2}_{x}\mathcal{V}^{1}\|_{L^{\infty}}$, $\|\partial^{2}_{x}\mathcal{V}^{2}\|_{L^{\infty}}$, respectively.
Step 1: Proof of (E1) for $\alpha>1$. Let $I(t)=\int_{\mathbb{R}}s_{\varepsilon}\,|x|^{2}\mathop{}\!\mathrm{d}x$. We sum up (1.3) and we multiply by $|x|^{2}$ to obtain
$\begin{split}&\partial_{t}\int_{\mathbb{R}}I(t)\!=\!-2\,\int_{\mathbb{R}}s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})\,x\mathop{}\!\mathrm{d}x-\!\int_{\mathbb{R}}(u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]+v_{\varepsilon}\,\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])x\mathop{}\!\mathrm{d}x+2\,\varepsilon\,M\\ &\leq\,2\,I(t)^{\frac{1}{2}}\,\left(\int_{\mathbb{R}}s_{\varepsilon}\,|\partial_{x}p(s_{\varepsilon})|^{2}\mathop{}\!\mathrm{d}x\right)^{\frac{1}{2}}+(\|\partial_{x}\mathcal{V}^{1}\|_{L^{\infty}}+\|\partial_{x}\mathcal{V}^{2}\|_{L^{\infty}})\,M^{\frac{1}{2}}\,I(t)^{\frac{1}{2}}+2\,\varepsilon\,M.\end{split}$ (3.2)
Similarly, multiplying by $p(s_{\varepsilon})$ and using Cauchy-Schwarz inequality we obtain
$\frac{\alpha}{\alpha-1}\partial_{t}\int_{\mathbb{R}}s_{\varepsilon}^{\alpha}\mathop{}\!\mathrm{d}x+\int_{\mathbb{R}}s_{\varepsilon}\,|\partial_{x}p(s_{\varepsilon})|^{2}\mathop{}\!\mathrm{d}x=-\!\int_{\mathbb{R}}(u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]+v_{\varepsilon}\,\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])\partial_{x}p(s_{\varepsilon})\mathop{}\!\mathrm{d}x\\ \leq\frac{1}{2}(\|\partial_{x}\mathcal{V}^{1}\|_{L^{\infty}}+\|\partial_{x}\mathcal{V}^{2}\|_{L^{\infty}})\,M+\frac{1}{2}\,\int_{\mathbb{R}}s_{\varepsilon}\,|\partial_{x}p(s_{\varepsilon})|^{2}\mathop{}\!\mathrm{d}x.$
Summing up the two identities and using Cauchy-Schwarz with a small parameter for the term $I(t)^{\frac{1}{2}}\,\left(\int_{\mathbb{R}}s_{\varepsilon}\,|\partial_{x}p(s_{\varepsilon})|^{2}\mathop{}\!\mathrm{d}x\right)^{\frac{1}{2}}$, we obtain the bound on $I(t)$ by the Grönwall’s inequality.
Step 2: Proof of (E1) for $\alpha\in(\frac{1}{3},1]$. For $\alpha\in(\frac{1}{3},1]$, we have
$\int_{\mathbb{R}}s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})\,x\mathop{}\!\mathrm{d}x=-\frac{1}{\alpha}\int_{\mathbb{R}}s_{\varepsilon}^{\alpha}\mathop{}\!\mathrm{d}x,$
so that computing as in (3.2) we have
$\partial_{t}\int_{\mathbb{R}}I(t)\!\leq\frac{2}{\alpha}\int_{\mathbb{R}}s_{\varepsilon}^{\alpha}\mathop{}\!\mathrm{d}x+(\|\partial_{x}\mathcal{V}^{1}\|_{L^{\infty}}+\|\partial_{x}\mathcal{V}^{2}\|_{L^{\infty}})\,M^{\frac{1}{2}}\,I(t)^{\frac{1}{2}}+2\,\varepsilon\,M.$
By splitting the set of integration for $|x|\leq 1$, $|x|>1$ and applying Hölder’s inequality we have
$\int_{\mathbb{R}}s_{\varepsilon}^{\alpha}\mathop{}\!\mathrm{d}x\leq\left(\int_{|x|\leq 1}s_{\varepsilon}\mathop{}\!\mathrm{d}x\right)^{\alpha}\,2^{1-\alpha}+\left(\int_{|x|>1}s_{\varepsilon}|x|^{2}\mathop{}\!\mathrm{d}x\right)^{\alpha}\,\left(\int_{|x|>1}|x|^{-\frac{2\alpha}{1-\alpha}}\mathop{}\!\mathrm{d}x\right)^{1-\alpha}.$
Since $\frac{2\alpha}{1-\alpha}>1$ for $\alpha>\frac{1}{3}$, we deduce $\int_{\mathbb{R}}s_{\varepsilon}^{\alpha}\mathop{}\!\mathrm{d}x\leq C\,(1+I(t))$ for a uniform constant $C$ and we conclude the proof of (E1) by the Grönwall’s inequality.
@@
Step 3: Proof of (E2), (E7) for $\alpha>\frac{1}{3}$. We multiply equations in (1.3) by $\log u_{\varepsilon}$ and $\log v_{\varepsilon}$, respectively, and sum them up to obtain
@@
$\begin{split}&\partial_{t}\int_{\mathbb{R}}u_{\varepsilon}\log u_{\varepsilon}\mathop{}\!\mathrm{d}x+\partial_{t}\int_{\mathbb{R}}v_{\varepsilon}\log v_{\varepsilon}\mathop{}\!\mathrm{d}x+\int_{\mathbb{R}}p^{\prime}(s_{\varepsilon})\,|\partial_{x}s_{\varepsilon}|^{2}\mathop{}\!\mathrm{d}x+\varepsilon\,\int_{\mathbb{R}}\frac{|\partial_{x}u_{\varepsilon}|^{2}}{u_{\varepsilon}}\mathop{}\!\mathrm{d}x+\varepsilon\,\int_{\mathbb{R}}\frac{|\partial_{x}v_{\varepsilon}|^{2}}{v_{\varepsilon}}\mathop{}\!\mathrm{d}x\\ &=-\int_{\mathbb{R}}\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\,\partial_{x}u_{\varepsilon}\mathop{}\!\mathrm{d}x-\int_{\mathbb{R}}\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]\,\partial_{x}v_{\varepsilon}\mathop{}\!\mathrm{d}x\leq(\|\partial_{x}^{2}\mathcal{V}^{1}\|_{L^{\infty}}+\|\partial_{x}^{2}\mathcal{V}^{2}\|_{L^{\infty}})\,M.\end{split}$ (3.3)
We note that $p^{\prime}(s_{\varepsilon})^{\frac{1}{2}}\,\partial_{x}s_{\varepsilon}=s_{\varepsilon}^{\frac{\alpha}{2}-1}\,\partial_{x}s_{\varepsilon}=\frac{2}{\alpha}\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}$. Hence, integrating the identity above in time and adding integrals of $\int_{\mathbb{R}}u_{\varepsilon}\,|x|^{2}\mathop{}\!\mathrm{d}x$, $\int_{\mathbb{R}}v_{\varepsilon}\,|x|^{2}\mathop{}\!\mathrm{d}x$ (which are controlled by (E1)) in order to control the negative part of the logarithms by means of Lemma 3.3, we conclude the proof of (E2) and (E7).
Step 4: Proof of (E1), (E2) and (E7) for $\alpha\in(0,\frac{1}{3}]$. In this range of $\alpha$, we will prove these assertions simultanously. The computation from Step 3 is still valid and we will rewrite it in the form
$\partial_{t}\int_{\mathbb{R}}u_{\varepsilon}\log u_{\varepsilon}\mathop{}\!\mathrm{d}x+\partial_{t}\int_{\mathbb{R}}v_{\varepsilon}\log v_{\varepsilon}\mathop{}\!\mathrm{d}x+\frac{4}{\alpha^{2}}\|\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}\|_{L^{2}_{x}}^{2}+\varepsilon\,\int_{\mathbb{R}}\frac{|\partial_{x}u_{\varepsilon}|^{2}}{u_{\varepsilon}}\mathop{}\!\mathrm{d}x+\varepsilon\,\int_{\mathbb{R}}\frac{|\partial_{x}v_{\varepsilon}|^{2}}{v_{\varepsilon}}\mathop{}\!\mathrm{d}x\leq C.$ (3.4)
Next, we sum up (1.3) and we multiply by $|x|$ to obtain
@@
We now study each of the terms appearing on the (RHS). Note that $s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})=\frac{1}{\alpha}\,\partial_{x}s_{\varepsilon}^{\alpha}$ so that splitting the integral for $x>0$ and $x<0$
@@
Note that $\alpha\leq\frac{1}{3}$ so that $\alpha<1-\alpha$ and we can estimate
$\frac{2}{\alpha}s_{\varepsilon}^{\alpha}(t,0)\leq\frac{2}{\alpha}+\frac{2}{\alpha}s_{\varepsilon}^{1-\alpha}(t,0)\leq\frac{2}{\alpha}+\frac{2}{\alpha}\int_{\mathbb{R}}|\partial_{x}s_{\varepsilon}^{1-\alpha}|\mathop{}\!\mathrm{d}x.$
Hence, integrating in time and using Lemma 3.2
$\begin{split}-\!\int_{0}^{t}\int_{\mathbb{R}}s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})\,\frac{x}{|x|}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}\tau&\leq\!\frac{2\,t}{\alpha}\!+\!\frac{2}{\alpha}\int_{0}^{t}\int_{\mathbb{R}}|\partial_{x}s_{\varepsilon}^{1-\alpha}|\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}\tau\leq\!\frac{2\,t}{\alpha}\!+\!C\,\|\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}\|_{L^{2}((0,t)\times\mathbb{R})}^{1+\frac{1-3\alpha}{1+\alpha}}\\ &\leq C\left(1+\|\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}\|_{L^{2}((0,t)\times\mathbb{R})}^{1+\frac{1-3\alpha}{1+\alpha}}\right)\end{split}$ (3.6)
by possibly modifying the constant $C=C(\alpha,T,M)$. Similarly,
$\begin{split}-\varepsilon\int_{\mathbb{R}}\partial_{x}s_{\varepsilon}\,\frac{x}{|x|}\mathop{}\!\mathrm{d}x&=2\,\varepsilon\,s_{\varepsilon}(t,0)=2\,\varepsilon\int_{0}^{\infty}|\partial_{x}u_{\varepsilon}|\mathop{}\!\mathrm{d}x+2\,\varepsilon\int_{0}^{\infty}|\partial_{x}v_{\varepsilon}|\mathop{}\!\mathrm{d}x\\ &\leq\frac{\varepsilon}{2}\int_{\mathbb{R}}\frac{|\partial_{x}u_{\varepsilon}|^{2}}{u_{\varepsilon}}\mathop{}\!\mathrm{d}x+\frac{\varepsilon}{2}\int_{\mathbb{R}}\frac{|\partial_{x}v_{\varepsilon}|^{2}}{v_{\varepsilon}}\mathop{}\!\mathrm{d}x+2\,\varepsilon\,M.\end{split}$ (3.7)
Finally, for the term with velocity fields
$\int_{\mathbb{R}}(u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]+v_{\varepsilon}\,\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])\,\frac{x}{|x|}\mathop{}\!\mathrm{d}x\leq(\|\partial_{x}\mathcal{V}^{1}\|_{L^{\infty}}+\|\partial_{x}\mathcal{V}^{2}\|_{L^{\infty}})\,M.$ (3.8)
Integrating (3.5) in time and plugging (3.6), (3.7) and (3.8) yields
$\int_{\mathbb{R}}s_{\varepsilon}(t,x)\,|x|\mathop{}\!\mathrm{d}x\leq$ $\int_{\mathbb{R}}s^{0}\,|x|\mathop{}\!\mathrm{d}x+C\left(1+\|\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}\|_{L^{2}((0,t)\times\mathbb{R})}^{1+\frac{1-3\alpha}{1+\alpha}}\right)$
$+\frac{\varepsilon}{2}\int_{0}^{t}\int_{\mathbb{R}}\left(\frac{|\partial_{x}u_{\varepsilon}|^{2}}{u_{\varepsilon}}+\frac{|\partial_{x}v_{\varepsilon}|^{2}}{v_{\varepsilon}}\right)\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}\tau$
for some constant $C=C(\alpha,T,M)$. We now observe that $1+\frac{1-3\alpha}{1+\alpha}<2$ so by Young’s inequality
$\|\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}\|_{L^{2}((0,t)\times\mathbb{R})}^{1+\frac{1-3\alpha}{1+\alpha}}\leq C(\delta,\alpha)+\delta\,\|\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}\|_{L^{2}((0,t)\times\mathbb{R})}^{2}$
for all $\delta\in(0,1)$. Hence, choosing $\delta$ appropriately, we obtain
@@
Step 5. Proof of (E4). Let $M({\gamma})=\sup_{t\in[0,T]}\left(\int_{\mathbb{R}}s_{\varepsilon}^{\gamma}\mathop{}\!\mathrm{d}x\right)^{\frac{1}{\gamma}}$. We will first prove that, for some constant $C=C(M,\alpha,T)$ independent of $\gamma$ and all $\gamma\geq 15\alpha+2$ and all $t\in[0,T]$
@@
We claim that (3.12) implies that $\{s_{\varepsilon}\}$ is bounded in $L^{\infty}(0,T;L^{q}(\mathbb{R}))$ for all $q\in[1,\infty)$ (although the argument will use only $q=3$). We split the proof for two cases.

### 16 Apr 2026 v2

@@ 3.1. A priori estimates
First, we obtain a priori estimates for the solutions of (1.4). While they are quite standard compared to the rest of the paper, they seem to be new in the fast-diffusion regime (particularly, the $L^{\infty}$ estimates). Moreover, one has to be careful as we work in the full range of exponents $\alpha\in(0,\infty)$ so for instance, when $\alpha\in(0,\frac{1}{3}]$, it is well-known that the energy $\int_{\mathbb{R}}s_{\varepsilon}^{\alpha}\mathop{}\!\mathrm{d}x$ is no longer integrable, nor is $\int_{\mathbb{R}}s_{\varepsilon}\,|x|^{2}\mathop{}\!\mathrm{d}x$, which can be deduced by looking at the Barenblatt profile, see e.g. [17, eq. (2.4)].
@@
Let $u_{\varepsilon}$, $v_{\varepsilon}$ be the solution of (1.4). Then, for all $T>0$, the following sequences are bounded:
@@
    $\{u_{\varepsilon}|x|\}$, $\{v_{\varepsilon}|x|\}$ in $L^{\infty}(0,T;L^{1}(\mathbb{R}))$,
@@
    $\{u_{\varepsilon}\log u_{\varepsilon}\,|x|^{\frac{1}{2}}\}$, $\{v_{\varepsilon}\log v_{\varepsilon}|x|^{\frac{1}{2}}\}$ in $L^{\infty}(0,T;L^{1}(\mathbb{R}))$ ,
@@
    $\{\partial_{t}u_{\varepsilon}\}$, $\{\partial_{t}v_{\varepsilon}\}$ in $L^{2}(0,T;H^{-1}(\mathbb{R}))$,
@@
We will need the following computation borrowed from [27, Prop. 2.1].
@@
$\int_{0}^{T}\int_{\mathbb{R}}s_{\varepsilon}^{2-\alpha}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}t\leq\|s^{0}\|_{L^{1}_{x}}\,\int_{0}^{T}\|s_{\varepsilon}^{1-\alpha}\|_{L^{\infty}(\mathbb{R})}\mathop{}\!\mathrm{d}t\leq\int_{0}^{T}\int_{\mathbb{R}}|\partial_{x}s_{\varepsilon}^{1-\alpha}|\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}t$
@@
We will also need the following lemma proved in [14, Lemma D.2].
@@
We divide the proof into a few steps. As a preliminary observation, we note that by the conservation of mass $M:=\int_{\mathbb{R}}s_{\varepsilon}(t,x)\mathop{}\!\mathrm{d}x=\int_{\mathbb{R}}s^{0}(x)\mathop{}\!\mathrm{d}x$, $\int_{\mathbb{R}}u_{\varepsilon}(t,x)\mathop{}\!\mathrm{d}x=\int_{\mathbb{R}}u^{0}(x)\mathop{}\!\mathrm{d}x$, $\int_{\mathbb{R}}v_{\varepsilon}(t,x)\mathop{}\!\mathrm{d}x=\int_{\mathbb{R}}v^{0}(x)\mathop{}\!\mathrm{d}x$. In particular, Assumption 1.1 implies that all the sequences $\{\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\}$, $\{\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]\}$, $\{\partial^{2}_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\}$, $\{\partial^{2}_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]\}$ are uniformly bounded in the space $L^{\infty}((0,T)\times\mathbb{R})$ and we denote the bound by $\|\partial_{x}\mathcal{V}^{1}\|_{L^{\infty}}$, $\|\partial_{x}\mathcal{V}^{2}\|_{L^{\infty}}$, $\|\partial^{2}_{x}\mathcal{V}^{1}\|_{L^{\infty}}$, $\|\partial^{2}_{x}\mathcal{V}^{2}\|_{L^{\infty}}$, respectively.
@@
Step 1: Proof of (E1), (E2) and (E7). We multiply equations in (1.4) by $\log u_{\varepsilon}$ and $\log v_{\varepsilon}$, respectively, and sum them up to obtain
@@
$\begin{split}&\partial_{t}\int_{\mathbb{R}}u_{\varepsilon}\log u_{\varepsilon}\mathop{}\!\mathrm{d}x+\partial_{t}\int_{\mathbb{R}}v_{\varepsilon}\log v_{\varepsilon}\mathop{}\!\mathrm{d}x+\int_{\mathbb{R}}p^{\prime}(s_{\varepsilon})\,|\partial_{x}s_{\varepsilon}|^{2}\mathop{}\!\mathrm{d}x+\varepsilon\,\int_{\mathbb{R}}\frac{|\partial_{x}u_{\varepsilon}|^{2}}{u_{\varepsilon}}\mathop{}\!\mathrm{d}x+\varepsilon\,\int_{\mathbb{R}}\frac{|\partial_{x}v_{\varepsilon}|^{2}}{v_{\varepsilon}}\mathop{}\!\mathrm{d}x\\ &=-\int_{\mathbb{R}}\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\,\partial_{x}u_{\varepsilon}\mathop{}\!\mathrm{d}x-\int_{\mathbb{R}}\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]\,\partial_{x}v_{\varepsilon}\mathop{}\!\mathrm{d}x\leq(\|\partial_{x}^{2}\mathcal{V}^{1}\|_{L^{\infty}}+\|\partial_{x}^{2}\mathcal{V}^{2}\|_{L^{\infty}})\,M.\end{split}$
We note that $p^{\prime}(s_{\varepsilon})^{\frac{1}{2}}\,\partial_{x}s_{\varepsilon}=s_{\varepsilon}^{\frac{\alpha}{2}-1}\,\partial_{x}s_{\varepsilon}=\frac{2}{\alpha}\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}$. Hence, we obtain
$\partial_{t}\int_{\mathbb{R}}u_{\varepsilon}\log u_{\varepsilon}\mathop{}\!\mathrm{d}x+\partial_{t}\int_{\mathbb{R}}v_{\varepsilon}\log v_{\varepsilon}\mathop{}\!\mathrm{d}x+\frac{4}{\alpha^{2}}\|\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}\|_{L^{2}_{x}}^{2}+\varepsilon\,\int_{\mathbb{R}}\frac{|\partial_{x}u_{\varepsilon}|^{2}}{u_{\varepsilon}}\mathop{}\!\mathrm{d}x+\varepsilon\,\int_{\mathbb{R}}\frac{|\partial_{x}v_{\varepsilon}|^{2}}{v_{\varepsilon}}\mathop{}\!\mathrm{d}x\leq C,$ (3.2)
for a constant $C$ independent of $\varepsilon$. Next, we sum up (1.4) and we multiply by $|x|$ to obtain
@@
We now consider the last two terms on the (RHS). For the first one
$\begin{split}-\varepsilon\int_{\mathbb{R}}\partial_{x}s_{\varepsilon}\,\frac{x}{|x|}\mathop{}\!\mathrm{d}x&=2\,\varepsilon\,s_{\varepsilon}(t,0)\leq 2\,\varepsilon\int_{0}^{\infty}|\partial_{x}u_{\varepsilon}|\mathop{}\!\mathrm{d}x+2\,\varepsilon\int_{0}^{\infty}|\partial_{x}v_{\varepsilon}|\mathop{}\!\mathrm{d}x\\ &\leq\frac{\varepsilon}{2}\int_{\mathbb{R}}\frac{|\partial_{x}u_{\varepsilon}|^{2}}{u_{\varepsilon}}\mathop{}\!\mathrm{d}x+\frac{\varepsilon}{2}\int_{\mathbb{R}}\frac{|\partial_{x}v_{\varepsilon}|^{2}}{v_{\varepsilon}}\mathop{}\!\mathrm{d}x+2\,\varepsilon\,M,\end{split}$
while for the term with the velocity fields
$\int_{\mathbb{R}}(u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]+v_{\varepsilon}\,\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])\,\frac{x}{|x|}\mathop{}\!\mathrm{d}x\leq(\|\partial_{x}\mathcal{V}^{1}\|_{L^{\infty}}+\|\partial_{x}\mathcal{V}^{2}\|_{L^{\infty}})\,M.$
Hence, from (3.3), we deduce
$\partial_{t}\int_{\mathbb{R}}s_{\varepsilon}\,|x|\mathop{}\!\mathrm{d}x\leq-\int_{\mathbb{R}}s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})\,\frac{x}{|x|}\mathop{}\!\mathrm{d}x+\frac{\varepsilon}{2}\int_{\mathbb{R}}\frac{|\partial_{x}u_{\varepsilon}|^{2}}{u_{\varepsilon}}\mathop{}\!\mathrm{d}x+\frac{\varepsilon}{2}\int_{\mathbb{R}}\frac{|\partial_{x}v_{\varepsilon}|^{2}}{v_{\varepsilon}}\mathop{}\!\mathrm{d}x+C,$ (3.4)
where $C$ does not depend on $\varepsilon\in(0,1)$. We will now combine (3.2) and (3.4) in order to conclude the proof. We need to split for two cases: $\alpha\leq 1$ and $\alpha>1$.
Step 1.1: $\alpha\leq 1$. Note that $s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})=\frac{1}{\alpha}\,\partial_{x}s_{\varepsilon}^{\alpha}$ so that we have
@@
Observe that $\alpha\leq 1$ implies $\alpha\leq\frac{1+\alpha}{2}$ so we can estimate
$\frac{2}{\alpha}s_{\varepsilon}^{\alpha}(t,0)\leq\frac{2}{\alpha}+\frac{2}{\alpha}s_{\varepsilon}^{\frac{1+\alpha}{2}}(t,0)\leq\frac{2}{\alpha}+\frac{2}{\alpha}\int_{\mathbb{R}}|\partial_{x}s_{\varepsilon}^{\frac{1+\alpha}{2}}|\mathop{}\!\mathrm{d}x\leq\frac{2}{\alpha}+\frac{2\,(1+\alpha)}{\alpha^{2}}\int_{\mathbb{R}}s_{\varepsilon}^{\frac{1}{2}}\,|\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}|\mathop{}\!\mathrm{d}x.$
Hence, integrating in time the estimates above and applying Cauchy-Schwarz inequality with a parameter $\delta$:
$\begin{split}-\!\int_{0}^{t}\int_{\mathbb{R}}s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})\,\frac{x}{|x|}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}\tau&\leq\frac{2\,t}{\alpha}+\frac{2\,(1+\alpha)}{\alpha^{2}}\int_{0}^{t}\int_{\mathbb{R}}s_{\varepsilon}^{\frac{1}{2}}\,|\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}|\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}\tau\\ &\leq C(t,\alpha,M,\delta)+\delta\,\|\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}\|_{L^{2}((0,t)\times\mathbb{R})}^{2}.\end{split}$
Plugging this estimate into (3.4), choosing $\delta=\frac{2}{\alpha^{2}}$, and integrating in time yields
@@
Step 1.2: $\alpha>1$. We sum up (1.4), we multiply by $p(s_{\varepsilon})$ and we use Cauchy-Schwarz inequality to obtain
$\begin{split}\frac{\alpha}{\alpha-1}\partial_{t}\int_{\mathbb{R}}s_{\varepsilon}^{\alpha}\mathop{}\!\mathrm{d}x+\int_{\mathbb{R}}s_{\varepsilon}\,|\partial_{x}p(s_{\varepsilon})|^{2}\mathop{}\!\mathrm{d}x=-\!\int_{\mathbb{R}}(u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]+v_{\varepsilon}\,\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])\partial_{x}p(s_{\varepsilon})\mathop{}\!\mathrm{d}x\\ \leq\frac{1}{2}(\|\partial_{x}\mathcal{V}^{1}\|^{2}_{L^{\infty}}+\|\partial_{x}\mathcal{V}^{2}\|^{2}_{L^{\infty}})\,M+\frac{1}{2}\,\int_{\mathbb{R}}s_{\varepsilon}\,|\partial_{x}p(s_{\varepsilon})|^{2}\mathop{}\!\mathrm{d}x.\end{split}$ (3.7)
It follows that $\{\sqrt{s_{\varepsilon}}\,\partial_{x}p(s_{\varepsilon})\}$ is bounded in $L^{2}((0,T)\times\mathbb{R})$. Combining with the bound on $\{\sqrt{s_{\varepsilon}}\}$ in $L^{2}((0,T)\times\mathbb{R})$, we deduce that the integral $\int_{0}^{T}\int_{\mathbb{R}}\left|s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})\,\frac{x}{|x|}\right|\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}\tau$ is uniformly bounded, so from (3.4) we obtain a bound on $\{s_{\varepsilon}\,|x|\}$ in $L^{\infty}(0,T;L^{1}(\mathbb{R}))$. The remaining assertions follow now from (3.2) together with Lemma 3.3, which allows us to control the negative part of the logarithm.
Step 2. Proof of (E4). Let $M({\gamma})=\sup_{t\in[0,T]}\left(\int_{\mathbb{R}}s_{\varepsilon}^{\gamma}\mathop{}\!\mathrm{d}x\right)^{\frac{1}{\gamma}}$. We will first prove that for some constant $C=C(M,\alpha,T)$ independent of $\gamma$ and all $\gamma\geq 15\,\alpha+2$, $t\in[0,T]$
@@
We claim that (3.9) implies that $\{s_{\varepsilon}\}$ is bounded in $L^{\infty}(0,T;L^{q}(\mathbb{R}))$ for all $q\in[1,\infty)$ (so that $M(\gamma)$ is finite for all $\gamma\geq 1$). We split the proof for two cases.